Torch alignment and wavelength calibration fails on 710 OES

Varian 710 oes, nebulization is ok, fog in the chamber is omogeneous,  tubing, nebulizer, torch, cone cleaned and alligned.

Work fine till 2 weeks ago, after i've experienced this: fail to execute wavelenght calibration end torch allign, No response for Mn 5ppm, No response for any element, fail to execute wavelenght calibration end torch allign.

I've noticed that stability of wavelenght calibration decreased during sequences in last period. (first standard alligned, Second standard after 10 samples slightly disalligned of  about + 0.1 nm), sometimes replies were not reproducible, 

Need help, technician is trying to understand what's the problem.

Many thanks


Parents Reply Children
No Data
Was this helpful?